OpenStudy (anonymous):

5^3sqrt(4x −1 + 2 = 17)

OpenStudy (anonymous):

\[5^3\sqrt4x −1 + 2 = 17\]

OpenStudy (anonymous):

is x under the square root?

OpenStudy (anonymous):

yes the sqrt goes to the 1

OpenStudy (anonymous):

1st subtract the 2 what do you get

OpenStudy (anonymous):

5^3sqrt(4x-1)=17-2

OpenStudy (anonymous):

5^3sqrt(4x-1)=17 125sqrt(4x-1)=17 sqrt(4x-1)=17/125 taking square 4x-1=(17/125)^2 4x=(17/125)^2+1 x=((17/125)^2+1)/4

OpenStudy (anonymous):

\[5^3\sqrt{4x-1}=15\] \[\sqrt{4x-1}=15/5^3\] \[(\sqrt{4x-1})^2=(15/5^3)^2\] \[4x - 1 = \frac{ 9 }{ 625 }\] \[4x = \frac{ 9 }{ 625 }+1\] \[4x = \frac{ 634 }{ 625 }\] \[x = \frac{ 317 }{ 1250}\]

OpenStudy (anonymous):

thank you guys :)

OpenStudy (anonymous):

he is wrong because 17-2= 15

OpenStudy (anonymous):

oops i made a mistake